May is building a fence around the outer edge of a rectangular garden plot that is 25ft by 45 ft. She will set the posts 5 feet apart. How many will she need?
Answer by checkley77(12844)

2*25+2*45=50+90=140 TOTAL FEET OF FENCING NEEDED.
140/5=28 POSTS WILL BE REQUIRED.